4(8-t)=20
6(2x-1)=4(2x+5)
7(2w-3)-5(3w-1)=0

Answer by rcmcc(152)

4(8-t)=20
6(2x-1)=4(2x+5)
7(2w-3)-5(3w-1)=0
I belive you are looking to solve for the variable in each case.
on the first.
4(8-t)=20
expand the brackets.
32-4t=20
subtract 32 from each side
-4t=-12
switch signs to make T positive
4t=12
divide out 4
t=3
next
6(2x-1)=4(2x+5)
expand brackets
12x-6=8x+20
put similar terms on different sides
12x-8x=20+6
4x=26
divide out 4.
x=6.5
the last one
7(2w-3)-5(3w-1)=0
expand the brackets
14w-21-15w+5=0
place similar terms on opposite sides
14w-15w=21-5
-w=18
switch signs to make w positive.
w=-18